OpenStudy (anonymous):

(x-2)(3x^2-6x+12) multiplying polynomials,,can anyone multiply and simplify?

OpenStudy (anonymous):

Use the distributive property. Distribute the x across all the terms in the second quantity and then distribute the -2 across. Combine like terms and you are done!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

(x+4) and (x-1) Why? Well, let us multiply them to see: (x+4)(x-1) = x(x-1) + 4(x-1) = x2 - x + 4x - 4 = x2 + 3x - 4 Multiplying (x+4)(x-1) together is called Expanding. In fact, Expanding and Factoring are opposites: Expanding is easy, but Factoring can often be tricky It can be like trying to find out what ingredients went into a cake to make it so delicious. It is sometimes not obvious at all! So let us try an example where we don't know the factors yet:

OpenStudy (anonymous):

Not factoring.

OpenStudy (anonymous):

who?

OpenStudy (anonymous):

I think he meant to type that in a different question?

OpenStudy (anonymous):

yea prob,that's not the same question i'll try to work it out